OpenStudy (anonymous):

Chase plays the piano and the cello. For every 2 hours he practices the piano, he practices the cello for 3 hours. If he practiced the piano for 7 hours last week, how many hours did he spend practicing the cello?

OpenStudy (anonymous):

@Michele_Laino @campbell_st

OpenStudy (campbell_st):

well this is a ratio question... so comparing the cello practice to the piano practice and keeping the same ratios... \[\frac{cello}{piano} = \frac{3}{2} = \frac{x}{7}\] just solve the equation for x

OpenStudy (anonymous):

how do i solve for x ?

OpenStudy (campbell_st):

multiply both sides of the equation by 7 and then calculate the answer